Find the ratio of the following:
81 to 108..................

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

81 / 108 = (3×3 × 3 × 3) / (2×2 × 3 × 3 × 3)
= 3/ 4

Thanks